OpenStudy (anonymous):

I'm not sure I did this right? Please help! I'm supposed to find the the tangent line to the ellipse

OpenStudy (anonymous):

I get a slope of 3/8...not - 3/8. Check your signs

OpenStudy (anonymous):

Checked them, and still have -3/8

OpenStudy (anonymous):

\[2x \frac{ dx }{ dx }+8y \frac{ dy }{ dx }=0\Rightarrow \frac{ dy }{ dx }=-\frac{ 2x }{8y}=-\frac{ x }{ 4y }\] slope at (3., -2) = 3/8

OpenStudy (anonymous):

Differentiating your original equation, did you get 2x + 8yy' = 0 ?

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

you have a negative slope when it should be positive

OpenStudy (anonymous):

ok,solve that for y', you get -2x / 8y which simplifies to -x / 4y

OpenStudy (anonymous):

Oh, I randomly put a negative there. That makes sense now. Thanks haha

OpenStudy (anonymous):

unbelievable